3 sin (x) + 4 cos (x) = 5. Find the value of sin (x). I have used (x) as theta

Grade:10

1 Answers

Vikas TU
7 years ago
Write  3 sin (x) + 4 cos (x) = 5
as:
3sinx = 5 – 4cosx
squaring both sides,
9(sinx)^2 = 25 + 16(cosx)^2 – 40cosx
9(1-(cosx)^2) = 25 + 16(cosx)^2 – 40cosx
9 – 25 = 25 (cosx)^2 – 40cosx
assume cosx = t and solve the quadratic.
25t^2 – 40t + 16 = 0
after then u can convert it into sinx.